OpenStudy (anonymous):

Jeff is baking a cake. The recipe says that he has to mix 32 grams of vanilla powder to the flour. Jeff knows that 1 cup of that particular vanilla powder has a mass of 128 grams. He added two over three of a cup of vanilla powder to the flour. Should Jeff add more vanilla powder to make the exact recipe or did he go over and by what amount? He needs to add one over four of a cup He went over by five over twelve of a cup He went over by one over four of a cup He needs to add five over twelve of a cup

OpenStudy (anonymous):



OpenStudy (anonymous):

we know he went over so it has to be B or C

OpenStudy (anonymous):

128 * 5/12 = 53.333 128* 1/4 = 32

OpenStudy (anonymous):

but we want 32 so 85 - 53 = 32